# Evaluating the Impact of Inter-cluster Communications in Edge Computing

### v1

@@ V-A Remote Maintenance (ML Inference)
Fig. 3 shows the response times achieved via the respective inter-cluster communication tools for ML inference in the Remote Maintenance use case, with the highest amount of payload that we tested of 1MB. In this case study, the sent data is echoed back to the sender, and we can observe the response time increasing exponentially with the worsening connection parameters between the nodes and clusters. The median value of the results at first exceeds our set threshold of 1000 ms with the edg20 parameter set for ClusterLink. The upper bound of the distribution however already exceeds this threshold with the edg15 parameter set, with the edg10 set barely remaining within the tolerable limit of 1000 ms. A higher round trip time than 15 ms between nodes is therefore not suitable to consistently maintain the requirements of our use case, assuming matching values for jitter and packet loss. For Submariner however, the upper bounds only slightly exceed the limits in the edg20 case, with the median value remaining around 300ms; significantly below critical values. Skupper exhibits slightly worse behavior than Submariner here, with its median value being almost double the one of Submariner in the edg20 case, but still within the boundary.
@@
As it can be seen by the increase from the smaller payload of 100KB in Fig. 4, a further increase of the payload size beyond 1MB would also result in less tolerance for connection parameters and therefore introduce stricter inter-node latency requirements to remain beneath the threshold, with all tools scaling in a similar fashion relative to the payload size.
@@
V-B Remote Maintenance (ML Training)
Fig. 5 shows the results of the ML training case studies for 1MB payload size which, differently from the inference use case, does not consider communication back to the request source besides acknowledgement. In contrast to the previous use case, we now see that ClusterLink performs better than Submariner, while Skupper now outperforms both; a trend that becomes more significant with the decreasing quality of inter-node networking conditions. Again, we can further see a significant jump in response times when crossing over from edg5 to edg10 scenario, pronouncing the exponential trend. Submariner and Clusterlink barely break the test limitations in the edg10 scenario with their upper bounds while Skupper manages to remain below critical values right until edg20 scenario. The median values cross our breaking point of 1000 ms at edg15 for Submariner and edg20 for ClusterLink, while at that point, Skupper’s median value still remains around 600 ms.
For smaller payloads, we observe in Fig. 6 that with 60 Bytes, Submariner provides the lowest response times, closely followed by Skupper, while ClusterLink falls behind them.
@@ V-C Vehicle Decision Assist
Fig. 7 shows the results for vehicle decision assist class of applications. It shows lower performance for ClusterLink compared to other tools, while Submariner outperforms Skupper and ClusterLink. Especially compared to Submariner, we can see a performance deterioration in ClusterLink for receiving data, which it does not exhibit when sending. The response times of all three tools however remain inside the set boundary of 100 ms, making each of them a suitable choice for this class of applications.
@@
When handling small payloads, the inter-node connection quality seems to only have linear influence on the total response time. We can confirm these findings by taking a look at the same payload size being transmitted both ways from test 1 as can be seen in Fig. 8. A further increase of the distance between the nodes beyond a round trip time of 20 ms would be unrealistic, since a higher response time for devices that are in danger of physically touching each other is highly unlikely.
Figure 8: Values from inference 1KB
@@ V-D7 Submariner Network Driver
We only explored Submariner’s VXLAN driver, while other research[11] has shown that its performance could vary when different drivers are used.
@@ VI Conclusion
In answering the research questions stated upfront, we proved that the response times have a linear relation to the geographical distance between all nodes when the transmitted data is low in size, while it can approach exponential growth with large payload sizes. We also observed that the combined impact of geographical spreading of nodes and payload size on the considered applications leads to exponential growth of application response times, at least for data being transferred into both directions and for data being sent. For the three tools analyzed, we found that ClusterLink outperformed Submariner only when “pushing” larger data to a service, was behind Submariner in all other cases and consistently outperformed by Skupper in all our tests. Submariner performed best when handling small amounts of data in all cases. Under these circumstances, it also exhibits the best behavior when the inter-node network parameters deteriorate. Finally, Skupper performed best when pushing larger data to a service but was slightly beaten by Submariner in all other cases.
It should be noted that the individual choice of software for a real-world setting should however always evaluate not just the performance, but also the feature set of the tools. For instance, if strong security is needed, ClusterLink’s authentication and access control might prevail, just as Submariner’s flexibility for connection establishment might make it more a more attractive choice. Also, when performance for bidirectional transfer of large data is the main concern, Skupper might be chosen as the best solution.

### v2

@@ V-A Remote Maintenance (ML Inference)
Fig. 3 shows the response times achieved via the respective inter-cluster communication tools for the ML inference in the Remote Maintenance use case with a payload of 1MB, the largest one in our value set. In this case study, the data sent is echoed back to the sender, generating a response of the same size as the request. We can observe a continuous, exponential increase of response times relative to the deteriorating networking conditions when using Submariner and Skupper, with Skupper achieving roughly two thirds of the response time of Submariner in our worst case. While Clusterlink is subject to the same trend, it seems to better deal with the additional delay imposed on the network connection between clusters, showing response times almost 50% lower than Skupper when using our lowest-quality links. In general, all tools were able to handle the request and response without their median response times exceeding the 1000 ms limit for the use case as defined in Table II. Submariner however occasionally exceeds the tolerable upper bound slightly when the parameter set with 20 ms of delay is introduced to the connection, which indicates that the tool is more sensible to link delay, delay variance and packet loss when compared to other tools.
@@
For smaller payloads, ClusterLink and Submariner performed similarly until the eg10 value set, with Submariner being slightly outperformed by ClusterLink in most cases as depicted in Fig. 4. However, the difference is subtle and in consideration of run to run variance, their performance can be considered equal here. Skupper, while initially delivering comparable performance to the other tools breaks away in the edg15 scenario by delivering ever higher response times, by almost 20% in the edg20 scenario, indicating a downside of using AMQP in this environment. From a response time perspective, this makes Skupper the last choice out of the three tools for inter-cluster communication when dealing with applications that need to send medium-sized payloads to a service that in turn replies with small payloads.
Figure 5: Results for Remote Maintenance (inference) edg15
In Fig. 5, we analyze the payload size’s impact on the tools’ performance for the remote maintenance inference use case in the scenario with the worst conditions that still allow all tools to perform below the boundary; edg15. We can observe similar performance for the three tools, with mean values below 100 ms for payloads up to 100 KB. With the payload size increased to 1MB, we see how the performance of the tools diverges, with drastic increases of the response times of Submariner, Skupper scaling slightly better, while ClusterLink’s jump in response times is significantly less pronounced.
We can therefore conclude that for workloads where small amounts of data are sent to remote services that reply with a similar amounts of data, the three tools perform similarly to each other. For medium and larger data, i.e. 100 KB and more, the tool selection becomes a decision of significant impact.
V-B Remote Maintenance (ML Training)
Fig. 6 shows the results of the ML training case studies for 1MB payload size, which, differently from the inference use case, does not consider communication back to the request source besides acknowledgments. Again, we can observe ClusterLink and Skupper behaving similarly to each other with slightly lower medians for ClusterLink, while it now also exhibits a slightly higher spread. Similarly to the ML inference for the remote maintenance use case with payloads of 1 MB, Submariner performs worst among the tested tools. Here, it presents about four times the median values of ClusterLink and Skupper in the edg20 scenario along with a high standard deviation in the observed data. In contrast to the previous use case, Submariner exceeds the boundary of 1000 ms already in the edg10 scenario with some measurements and in the edg15 scenario with its median value. ClusterLink and Skupper however remain below the threshold with their median values but slightly exceed the boundary in the edg20 scenario with a small portion of the observed samples.
@@ V-B Remote Maintenance (ML Training)
The results for experiments with smaller payloads of 60 Bytes are presented in Fig. 7. The plots show almost identical performance for the three inter-cluster communication tools, to the level of some even reaching identical median values in the edg15 and edg20 scenario.
@@
In Fig. 8, we show the impact of varying payload sizes on the tools’ performance for the remote maintenance training use case under the worst networking conditions that still allow all tools to achieve median response times below the boundaries. We can observe similar performance for all three tools until the payload size exceeds 2 KB. Higher payloads lead to Submariner’s performance degrading significantly quicker than for ClusterLink and Skupper in relation to the payload size, indicating a clear advantage of the mTLS-based implementations over the IPsec-based Submariner due to less overhead of this approach. ClusterLink then brings this overhead down further by utilizing HTTP instead of AMQP for transport, resulting in even better performance than Skupper.
Curiously, the worst performance of Submariner that we observed was when data was not sent back to the tester container (or client), apart from acknowledgments. We can therefore conclude, that Submariner’s implementation seems to have suboptimal handling of HTTP POST responses when it comes to larger payloads, which makes it unsuitable for this application when network latency is above 7.5 ms in each direction (assuming a symmetric link, therefore 15 ms for round trip).
Figure 8: Results for Remote Maintenance (Training) edg15
@@ V-C Vehicle Decision Assist
Fig. 9 shows the results for the vehicle decision assist class of applications. ClusterLink and Skupper again present very similar performance with their median values being identical across almost all scenarios. Submariner however seems to consistently perform slightly better (1 ms lower median response times) than the other tools. When handling these small payloads, according to our results, the inter-node connection quality only has a linear influence on the total response time. The response times of all three tools remain inside the set boundary of 100 ms, making each of them a suitable choice for this class of application. A further increase of the distance between the nodes beyond a round trip time of 20 ms would be unrealistic, since such a high latency for devices that are in danger of physically touching each other is highly unlikely.
@@ V-D7 Submariner Network Driver
We only explored Submariner’s libreswan (IPsec) driver, while other research [11] has shown that its performance could vary when different drivers are used.
@@ VI Conclusion
In answering the research questions stated upfront, we proved that the response times have a linear relation to the geographical distance between all nodes when the transmitted data is low in size, while it can approach exponential growth with large payload sizes. We also observed that the combined impact of geographical spreading of nodes and increasing payload size on the considered applications leads to exponential growth of application response times, at least for data being transferred into both directions and for data being sent. For the three tools analyzed, we found that ClusterLink outperformed Submariner and Skupper whenever larger data was exchanged with a service, Submariner performed slightly better for small payloads that are received via HTTP GET requests and Skupper performed decently unless medium-sized payloads had to be pushed via HTTP POST requests.
In more general terms, we can say that for inter-cluster communication, creating network layer tunnels like Submariner seems to provide higher efficiency for small payloads but suffers when dealing with large amounts of data as soon as significant delay, variance and packet loss is experienced on the connecting link. Solutions with the ability of proxying TCP connections are at advantage here, delivering lower overall response times. Furthermore, use cases where data has to be sent to another service with significantly sized replies seem to suffer when using Skupper or Submariner in edge networks with suboptimal inter-node connections.
It should be noted that the individual choice of software for a real-world setting should however always evaluate not just the performance, but also the feature set of the tools. For instance, if strong security is needed, ClusterLink’s authentication and access control might prevail, just as Submariner’s flexibility for connection establishment might make it more a more attractive choice. Also, when simplicity of the configuration setup is the main concern, Skupper might be chosen as the simplest solution.